DQ Industries is analyzing their capital budgeting options and need to make sure their WACC is correct. The current T-Bill rate is 2.5%, market return 10%, and their beta is .9. Their capital structure consists of $6,000,000 in equity and $4,000,000 in 20 year bonds that sold 7 years ago for $1010 but $20 in flotation costs were paid. The coupon rate is 5.4% and payment is semi-annual. Tax rate is 40%.

What is their cost of debt?

What is their cost of equity?

What is their WACC?

Continuing with DQ Industries, now that they have their current WACC, if they go forwards with the proposed capital project they will need an additional $2,000,000, which will be obtained via a loan at a stated interest rate of 5.5%. What will be the new WACC?
Answer #1
COST OF DEBT
Bond Face Value $1,000
Selling Price per bond $1,010
Flotation Cost per bond $20
Pv Net amount received per bond $990
Nper Number of semi annual periods                     40 (20*2)
Pmt Semi annual coupon payment $27.00 (1000*5.4%)/2
Fv Amount to be paid on maturity $1,000
RATE Semi annual yield to maturity 2.7415% (Using RATE function of excel with Nper=40, Pv=-990, Pmt=27, Fv=1000)
Before tax annual cost of debt 5.4829% (2.7415*2)
After Tax Cost of Debt =5.4829*(1-Tax Rate)
Cd After Tax Cost of Debt = 3.29% (5.4829*(1-0.4)
COST OF EQUITY
Rs=Rf+Beta*(Rm-Rf)
Rs=Required Return on shares
Rf=Risk free rate =2.5%
Rm=Market return=10%
Beta=0.9
Required Return=2.5+0.9*(10-2.5)= 9.25%
Ce Cost of Equity 9.25%
A Amount of Equity $6,000,000
B Amount of Debt $4,000,000
C Total Capital $10,000,000
We=A/C Weight of Equity                 0.60
Wd=B/C Weight of Debt                 0.40
WACC=Cd*Wd+Ce*We
WACC=3.29%*0.4+9.25%*0.6= 6.87%
Loan amount $2,000,000
Cl After tax cost of loan =5.5*(1-0.4) 3.30%
D Amount of Equity $6,000,000
E Amount of Debt $4,000,000
F Loan amount $2,000,000
G Total Capital $12,000,000
We=D/G Weight of Equity                 0.50
Wd=E/G Weight of Debt                 0.33
Wl=F/G Weight of Loan                 0.17
WACC=We*Ce+Wd*Cd+Wl*Cl
WACC=3.29%*0.33+9.25%*0.5+3.30%*0.17 6.27%
NEW WACC =6.27% D11 =RATE(D8,09,-D7,010) А в D E F G H I Pv COST OF DEBT Bond Face Value Selling Price per bond Flotation Cost per bond Net a